# Modular Compilation for Quantum Chiplet Architectures

### v2

@@ 8. Conclusion
Physical, technological and economic considerations make the prospect of a single “monolithic” large-scale quantum processor very expensive or entirely unrealistic by current standards. These constraints force future quantum systems to adopt a physically distributed architecture, based on modular quantum chiplets. As such, modular quantum processors already dominate the roadmaps of several major quantum companies.
However, the complexity of chiplet-based quantum devices, coupled with their growing size, presents an imminent scalability challenge for quantum compilation. Existing qubit allocation methods are often unable to contend with inter-chiplet links, which don’t necessary support a universal basis gate set, and existing methods of logical-to-physical qubit placement, routing, unitary synthesis, and quantum circuit optimization, are typically not designed for qubit links that vary significantly in terms of duration or fidelity.
In this work, we first modify stock contemporary compilers to become aware of hardware modularity, thereby allowing them to correctly compile circuits for modular architectures with limited cross-chiplet gate support. We then propose SEQC, a Stratify-Elaborate Quantum Compiler for modular architectures. SEQC performs compilation in two stages, with the first stage (stratification, or chiplet splitting) performed only once for a given architecture, and the second stage (elaboration, or chiplet compilation) performed in parallel for each chiplet. Unlike contemporary quantum compilers that have to compile a quantum circuit anew before each execution, only the elaboration stage of SEQC needs to be performed recurrently before each execution. As its computational complexity is a function of the quantum chiplet size, not the entire quantum processor size, and chiplet sizes grow slowly, SEQC’s recurrent compilation time is largely unaffected by the relentless growth of qubit counts in future quantum processors.
We design and implement several novel methods for qubit placement, qubit routing, and circuit optimization in SEQC, with both hardware modularity and compilation parallelization in mind. Owning to these techniques and its parallelization strategy, SEQC compiles circuits that exhibit up to a 36% increase in circuit fidelity and up to 1.92×1.92$\times$1.92 × lower execution time, while consistently achieving 24×2-4$\times$2 - 4 × faster compilation time over a chiplet-aware Qiskit baseline.

### v3

@@ 8. Conclusion
Physical, technological, and economic considerations force future quantum systems to adopt a physically distributed architecture. Modular quantum processors already dominate the roadmaps of several major quantum companies. However, quantum compilers face a scalability challenge brought by the complexity and increasing scale of chiplet-based quantum devices, and their lack of chiplet awareness impedes their ability to produce high-quality results.
In this work, we first modify stock contemporary compilers to become aware of hardware modularity, thereby allowing them to correctly compile circuits for modular architectures with limited cross-chiplet gate support. We then propose SEQC, a Stratify-Elaborate Quantum Compiler for modular architectures. SEQC performs compilation in two stages, with the first stage (stratification, or chiplet splitting) performed only once for a given architecture, and the second stage (elaboration, or chiplet compilation) performed in parallel for each chiplet. Unlike contemporary quantum compilers that have to compile a quantum circuit anew before each execution to capitalize on the latest qubit calibration results, only the elaboration stage of SEQC needs to run before each execution. As its computational complexity is a function of the quantum chiplet size, not the entire quantum processor size, and chiplet sizes grow slowly, SEQC’s recurrent compilation time is largely unaffected by the relentless growth of qubit counts in future quantum processors.
SEQC implements several novel methods for qubit placement, qubit routing, and circuit optimization, with both hardware modularity and compilation parallelization in mind, achieving on average $9.3\%$ higher circuit fidelity (up to $49.99\%$), while attaining on average 3.27×3.27$\times$3.27 × faster compilation (up to 6.74×6.74$\times$6.74 ×) compared to a chiplet-unaware Qiskit baseline.